# FoMEMO: Towards Foundation Models for Expensive Multi-objective Optimization

### v1

@@ E.4 Parallel Scalability
Table 7: Means (stds) of best hypervolume metrics ($\uparrow$) in parallel optimization (q=10) on real-world applications over 10 independent runs. The best and second best mean results for each row are highlighted with shaded background and underlines respectively.
Test Instances Sobol qParEGO qEHVI qNEHVI FoMEMO-EI FoMEMO-UCB FoMEMO-UHVI
RE21 7.217e-01(9.10e-03) 8.481e-01(1.05e-02) 8.813e-01(4.48e-04) 8.823e-01(1.35e-04) 8.647e-01(1.59e-03) 8.701e-01(1.39e-03) 8.592e-01(2.11e-03)
RE22 5.695e-01(2.01e-02) 5.272e-01(3.00e-02) 5.515e-01(3.69e-02) 5.852e-01(2.88e-02) 5.461e-01(4.45e-02) 5.422e-01(5.44e-02) 5.978e-01(4.80e-02)
RE23 3.216e-01(1.53e-01) 1.116e+00(1.05e-02) 1.134e+00(3.43e-03) 1.132e+00(6.00e-03) 1.152e+00(2.89e-03) 1.150e+00(3.06e-03) 9.437e-01(8.27e-02)
RE24 9.873e-01(4.37e-02) 1.166e+00(9.05e-04) 1.138e+00(6.54e-03) 1.149e+00(5.45e-03) 1.156e+00(3.84e-03) 1.156e+00(3.01e-03) 1.090e+00(2.08e-02)
RE25 2.011e-01(1.58e-01) 8.600e-01(7.47e-02) 1.018e+00(1.09e-02) 9.487e-01(2.99e-02) 9.626e-01(6.56e-02) 1.007e+00(2.71e-02) 7.345e-01(9.27e-02)
RE31 1.246e+00(2.97e-02) 1.312e+00(1.72e-02) 1.293e+00(9.40e-03) 1.304e+00(3.66e-03) 1.308e+00(5.46e-03) 1.309e+00(3.07e-03) 1.304e+00(1.91e-02)
RE32 1.271e+00(1.82e-02) 1.319e+00(4.92e-03) 1.324e+00(3.03e-03) 1.322e+00(1.90e-03) 1.323e+00(2.85e-03) 1.324e+00(2.00e-03) 1.293e+00(8.26e-03)
RE33 1.125e+00(3.16e-02) 1.273e+00(3.43e-02) 1.311e+00(6.10e-04) 1.311e+00(9.34e-04) 1.311e+00(9.51e-04) 1.312e+00(8.16e-04) 1.281e+00(9.48e-03)
RE34 5.787e-01(1.08e-02) 9.348e-01(1.69e-02) 1.036e+00(5.77e-04) 1.037e+00(8.88e-04) 9.920e-01(9.27e-03) 1.008e+00(8.27e-03) 9.308e-01(6.52e-03)
RE35 1.181e+00(1.56e-02) 1.272e+00(7.54e-03) 1.298e+00(7.60e-04) 1.297e+00(1.10e-03) 1.287e+00(6.37e-03) 1.288e+00(8.94e-03) 1.257e+00(8.02e-03)
RE36 3.233e-01(6.46e-02) 8.424e-01(3.12e-02) 8.912e-01(1.00e-02) 8.865e-01(6.89e-03) 9.016e-01(5.60e-03) 9.073e-01(4.82e-03) 8.502e-01(1.17e-02)
RE37 6.216e-01(2.30e-02) 7.996e-01(2.07e-02) 8.850e-01(1.10e-03) 8.865e-01(8.78e-04) 8.454e-01(5.26e-03) 8.591e-01(4.19e-03) 8.279e-01(6.17e-03)

### v2

@@ D.1 Dimensional Scalability
Table 9: Means (stds) of hypervolume metrics ($\uparrow$) on 3 engineering design problems with many objectives and a multi-modal WFG problem with complex search space over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm Car Side Impact (7d4m) Conceptual Marine (6d4m) Water Resource (3d6m) WFG (30d3m)
MOEA/D-EGO 7.039e-01(5.7e-03) 5.915e-01(2.4e-02) 1.412e+00(1.3e-02) 2.754e-01(5.3e-03)
TSEMO 6.942e-01(7.9e-03) 7.358e-01(7.9e-03) 1.198e+00(5.7e-02) 2.869e-01(4.3e-03)
USEMO-EI 7.151e-01(7.2e-03) 6.750e-01(1.1e-02) 1.433e+00(9.4e-03) 2.842e-01(5.3e-03)
DGEMO N/A N/A N/A 2.792e-01(4.6e-03)
qNEHVI 5.900e-01(7.0e-03) 6.472e-01(2.2e-02) 1.342e+00(4.8e-02) 3.276e-01(2.2e-02)
qParEGO 5.740e-01(2.3e-02) 3.887e-01(2.6e-02) 1.404e+00(3.7e-02) 3.213e-01(1.5e-02)
JES 5.629e-01(1.7e-02) 4.628e-01(2.9e-02) 1.296e+00(2.6e-02) 3.314e-01(7.5e-03)
BOFormer N/A N/A N/A 3.086e-01(5.2e-03)
FoMEMO-EI 7.538e-01(7.4e-03) 7.505e-01(1.9e-02) 1.470e+00(4.7e-03) 3.159e-01(4.3e-03)
FoMEMO-UCB 7.739e-01(8.1e-03) 7.814e-01(1.5e-02) 1.475e+00(3.9e-03) 3.150e-01(4.4e-03)
FoMEMO-UHVI 7.156e-01(4.7e-03) 6.012e-01(1.6e-02) 1.430e+00(1.9e-02) 3.173e-01(7.8e-03)
FoMEMO-UR2I 7.552e-01(5.1e-03) 7.587e-01(3.8e-03) 1.370e+00(2.0e-02) 3.397e-01(1.4e-02)
In this subsection, we further evaluate the dimensional scalability of our method, including both objective and feature dimensions. Firstly, we consider 3 more challenging real-world engineering design problems with higher objective dimensions, including car side impact design and conceptual marine design problems, both with 4 objectives, as well as water resource planning problem, which has 6 objectives (Tanabe and Ishibuchi, 2020). The experimental results are shown in Table 9. We find that our method demonstrates superior scalability and maintains robust performance in many-objective optimization scenarios. A primary factor contributing to this efficacy of our approach is its ability to directly predict the exact aggregated posterior as a universal basis, which avoids the performance-degrading compromises often found in existing methods. In contrast, traditional methods suffer from significant limitations as the objective dimensionality increases: qNEHVI faces the computational burden of non-dominated partitioning and hypervolume calculations, which become computationally intractable in high-dimensional spaces. MOEA/D-EGO introduces approximation errors due to the non-Gaussian nature of the max operation when using Gaussian distribution to estimate the unknown aggregation distribution in high-dimensional objective spaces. By bypassing these approximations, our method preserves modeling fidelity without sacrificing efficiency. Secondly, unlike qParEGO that model each sub-problem repeatedly, our method utilizes the advantage of similar patterns shared by sub-problems with adjacent preferences, which means the belief updates of each sub-problem during training or optimization can be efficiently facilitated by its neighbors.
To examine the scalability of our approach across varying feature dimensions, we employ the WFG4 benchmark (Huband et al., 2006), which allows for flexible expansion of the decision space. Characterized by its pronounced multi-modality, WFG4 presents a complex landscape that rigorously evaluates an algorithm’s ability to maintain global exploration and circumvent premature convergence. For this evaluation, we consider a 30-dimensional instance as a representative case. This setup aligns with the maximum dimensionality encountered during our current training phase, serving as a critical testbed to verify the model’s generalization robustness at the boundary of its training scope. Experimental results show that FoMEMO-UR2I consistently maintains superior performance even in this challenging scenario. While our architecture is readily adaptable to higher-dimensional spaces through expanded training, we focus on the aforementioned configurations here to establish a performance baseline for dimensionality scalability. Expanding the current framework to scale toward even more massive objective and decision spaces has been designated as a primary objective for our future work.
D.2 Parallel Scalability
In this subsection, we provide supplementary experimental results to further validate the parallel scalability of our approach. Tables 10 and 11 report the performance across all algorithms with batch sizes of $q=5$ and $q=10$ per iteration, respectively. To ensure a rigorous and consistent comparison, a total evaluation budget of 100 samples is allocated for each configuration, excluding the initial design points. Our findings reveal that even with a straightforward batch generation strategy, our method consistently outperforms methods such as TSEMO and DGEMO in most cases, despite their reliance on sophisticated batch selection heuristics. Notably, among our proposed variants, the two preference-based acquisition functions yield the most superior results. This performance gap underscores the advantage of leveraging a foundation model pre-trained on diverse synthetic landscapes. Furthermore, it demonstrates that modeling aggregated posteriors parameterized by preferences is inherently amenable to parallelization, providing a more robust and natural framework for multi-objective optimization without focusing on developing more sophisticated batch-selection strategies.
Table 10: Means (stds) of hypervolume metrics ($\uparrow$) in parallel optimization (q=5) on representative problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm BC ABD Four Bar Truss Gear Train Materials Mic
MOEA/D-EGO 7.052e-01(1.5e-02) 6.227e-01(5.3e-02) 7.658e-01(1.7e-02) 5.365e-01(1.2e-01) 1.164e+00(6.2e-03) 1.247e+00(9.0e-03)
TSEMO 8.081e-01(5.8e-04) 7.117e-01(1.5e-02) 8.520e-01(1.9e-03) 7.811e-01(3.8e-02) 1.155e+00(5.9e-03) 1.238e+00(7.5e-03)
USEMO-EI 7.882e-01(5.6e-03) 7.057e-01(1.8e-02) 8.287e-01(6.3e-03) 6.742e-01(4.9e-02) 1.150e+00(5.3e-03) 1.243e+00(1.5e-02)
DGEMO 8.068e-01(8.8e-04) 5.838e-01(5.9e-02) 8.549e-01(3.0e-03) 8.413e-01(6.8e-02) 1.162e+00(7.8e-03) 1.272e+00(1.8e-02)
qNEHVI 7.990e-01(1.5e-03) 7.105e-01(1.1e-02) 8.240e-01(5.8e-03) 7.810e-01(7.5e-02) 1.150e+00(7.9e-03) 1.232e+00(2.1e-02)
qParEGO 7.744e-01(3.3e-03) 6.852e-01(8.5e-03) 7.651e-01(1.8e-02) 6.832e-01(7.2e-02) 1.142e+00(1.6e-02) 1.176e+00(2.1e-02)
JES 7.575e-01(5.5e-03) 6.601e-01(1.6e-02) 7.861e-01(9.1e-03) 5.783e-01(7.0e-02) 1.157e+00(4.3e-03) 1.211e+00(2.1e-02)
FoMEMO-EI 8.064e-01(1.2e-03) 7.351e-01(8.0e-03) 8.656e-01(2.2e-03) 9.043e-01(7.1e-03) 1.171e+00(2.2e-03) 1.245e+00(1.4e-02)
FoMEMO-UCB 8.072e-01(5.3e-04) 7.311e-01(5.1e-03) 8.698e-01(1.5e-03) 9.229e-01(5.3e-03) 1.176e+00(2.4e-03) 1.236e+00(1.9e-02)
FoMEMO-UHVI 8.036e-01(6.4e-04) 7.298e-01(6.5e-03) 8.567e-01(1.7e-03) 8.381e-01(1.2e-02) 1.173e+00(1.5e-03) 1.250e+00(2.5e-02)
FoMEMO-UR2I 7.922e-01(1.8e-03) 6.538e-01(9.8e-03) 8.420e-01(4.5e-03) 8.288e-01(2.1e-02) 1.156e+00(4.5e-03) 1.214e+00(1.2e-02)
Table 11: Means (stds) of hypervolume metrics ($\uparrow$) in parallel optimization (q=10) on representative problems over 10 independent runs. The best, second-best, and third-best mean results for each problem are highlighted with dark gray, gray, and light gray backgrounds, respectively.
Algorithm BC ABD Four Bar Truss Gear Train Materials Mic
MOEA/D-EGO 7.221e-01(2.1e-02) 5.363e-01(6.9e-02) 7.872e-01(2.0e-02) 3.939e-01(1.3e-01) 1.154e+00(5.7e-03) 1.238e+00(1.1e-02)
TSEMO 8.074e-01(7.6e-04) 6.512e-01(3.3e-02) 8.521e-01(2.1e-03) 7.060e-01(6.9e-02) 1.145e+00(9.6e-03) 1.242e+00(9.8e-03)
USEMO-EI 7.708e-01(1.4e-02) 6.633e-01(2.2e-02) 8.267e-01(1.2e-02) 6.155e-01(9.7e-02) 1.143e+00(7.9e-03) 1.231e+00(1.4e-02)
DGEMO 8.063e-01(1.3e-03) 5.228e-01(5.8e-02) 8.713e-01(3.0e-03) 8.221e-01(5.7e-02) 1.149e+00(7.3e-03) 1.242e+00(1.3e-02)
qNEHVI 7.975e-01(2.4e-03) 7.222e-01(7.5e-03) 8.194e-01(3.9e-03) 7.600e-01(8.7e-02) 1.145e+00(9.2e-03) 1.207e+00(2.1e-02)
qParEGO 7.722e-01(7.8e-03) 6.892e-01(1.2e-02) 7.286e-01(2.3e-02) 6.207e-01(1.0e-01) 1.150e+00(1.1e-02) 1.175e+00(1.4e-02)
JES 7.622e-01(3.2e-03) 6.245e-01(2.1e-02) 7.813e-01(8.3e-03) 5.677e-01(6.8e-02) 1.159e+00(3.0e-03) 1.222e+00(1.6e-02)
FoMEMO-EI 8.063e-01(8.5e-04) 7.248e-01(7.4e-03) 8.652e-01(1.8e-03) 8.987e-01(1.5e-02) 1.164e+00(4.7e-03) 1.237e+00(2.7e-02)
FoMEMO-UCB 8.064e-01(6.5e-04) 7.347e-01(1.2e-02) 8.679e-01(2.8e-03) 9.188e-01(6.1e-03) 1.167e+00(5.4e-03) 1.247e+00(2.6e-02)
FoMEMO-UHVI 7.881e-01(2.1e-02) 7.179e-01(1.0e-02) 8.494e-01(1.7e-03) 8.385e-01(1.6e-02) 1.166e+00(4.8e-03) 1.259e+00(1.5e-02)
FoMEMO-UR2I 7.297e-01(2.5e-02) 5.862e-01(1.9e-02) 7.577e-01(1.7e-02) 6.736e-01(6.4e-02) 1.137e+00(1.2e-02) 1.180e+00(2.0e-02)
Appendix E Limitations and Potential Future Work
E.1 GP Priors Limitations
In this paper, we focus on addressing a general class of unconstrained multi-objective expensive optimization problems characterized by continuous, nonlinear objective functions that exhibit a degree of smoothness and allow for uncertainty quantification, which are prevalent in numerous real-world applications, making them well-suited for Gaussian Process (GP) modeling. Therefore, we can extensively sample from GP priors to generate a substantial amount of synthetic training data, simulating a wide range of potential mappings that may exist in such real-world scenarios. The experimental results further demonstrate the applicability and generalization of our method. However, when confronted with problems that involve special knowledge completely beyond GP priors, our method may struggle to generalize to these specific types of cases, as we cannot train the model indefinitely to traverse the infinite space of function hyperparameters. Potential solutions include incorporating expert knowledge to design specialized priors during synthetic data generation, or interacting with the evaluation environment during optimization to discover and enrich potential priors. Since synthetic data can be generated from the designed priors through extensive sampling, these enhancements can be seamlessly integrated into our existing framework. Furthermore, instead of permanently freezing the model parameters after a single round of synthetic pre-training, it may be advantageous to fine-tune the pre-trained model using a small amount of domain-specific data.
E.2 Scalability Challenges
Although we have evaluated the scalability of our approach across problems with varying feature and objective spaces (see Section D.1), those involving significantly higher-dimensional features or objectives pose considerable challenges. While our framework can easily accommodate higher-dimensional data during training, the sample complexity required to adequately represent such spaces scales exponentially, which imposes a prohibitive challenge to the synthesis of a representative pre-training corpus. This curse of dimensionality mirrors a fundamental bottleneck in the multi-objective optimization community, where the exponential expansion of the search space remains a primary challenge for both surrogate modeling and search efficiency. To this end, several promising avenues worth further exploration. In the context of high-dimensional feature spaces, the development of efficient dimensionality reduction methods and embedding techniques, along with effective search strategies such as the trust region Bayesian optimization (TuRBO) algorithm (Eriksson et al., 2019), presents significant potential. For problems involving a large number of objectives, the number of solutions required to approximate the entire Pareto set increases exponentially, resulting in significant overhead for both optimization and decision-making. Therefore, rather than pursuing a comprehensive set of Pareto solutions, recent work has focused on identifying a small set of representative solutions, ensuring that each objective can be well addressed by at least one solution within the set (Lin et al., 2024a; Liu et al., 2024). This motivates the development of foundational models and parallel acquisition functions based on the set scalarizations (Lin et al., 2024a).